# OSS Malicious Package Analysis in the Wild

### v2

@@ III-A Node & Edge
Co-existing Edge indicates that two malicious packages co-exist in the same security report. As we mentioned before, the security report is the evidence that reveals the attack campaign behind the malicious package. A security report may contain multiple malicious packages, indicating a multi-to-one mapping between a report and packages. For example, the security report [58] claimed that an adversary author ‘Lolip0p’ developed malicious packages called ‘Colorslib’, ‘httpslib’ and ‘libhttps’, where the co-existing relationship ‘Colorslib’, ‘httpslib’, ‘Colorslib’, ‘libhttps’, and ‘httpslib’, ‘libhttps’.
@@
A dataset with the security report is essential to build the co-existing edge in the malicious package graph. However, there is no public dataset with the security report. To address this practical challenge, we propose a web crawler to collect the security analysis reports from the Internet. (1) We use the website URL list (Section II-B) as the initial seed for finding the security analysis reports. (2) We leverage the search engine (Google) to find similar URLs to the initial URL. (3) We use the web crawler to collect the security analysis reports from the Internet. (4) We manually filter out irrelevant web pages and keep the security reports about malicious packages. Through the manual approach, we have found 68 websites that provide security analysis reports about malicious packages. Table III lists the online sources of websites, and we have found 1,366 security reports. Our dataset is the largest number of security reports about malicious packages in the OSS ecosystem.
@@
III-B Connected Subgraph
We extract connected subgraphs to represent malicious package groups based on the four types of edges. Table II lists the detailed information of the subgraph, including the node number, the edge number, the out-degree, and the in-degree.
Duplicated Group (DG) characterizes the same malicious package from different sources, indicating their overlapping degree. Malicious package detection is conducted independently, where the cooperation of different research organizations seems rare. If two nodes have a duplicated edge , we put them () into the same DG.
Dependency Group (DeG) characterizes the collective set where each malicious package has a dependency relationship with others. Typically, an SSC attack campaign employs multiple malicious packages, where the first package, pretending to be innocent, uses a dependency malicious library to hide its malicious behavior. When a package uses a different malicious package as its dependency library, it is highly probable that they were developed by the same attacker or criminal. If two nodes have a dependency edge , we put them into the same DeG.
Similarity group (SG) represents the collective set where malicious packages have similar source codes. Adversaries cannot use the same package name after the package manager removes it. Hence, adversaries must change the package to release a new malicious package with a different name for the subsequent attack. The same adversary may release several malicious packages over a long period, where those packages may use identical code bases. If two nodes have a similar edge , we put them into the same SmG.
Co-existing Group (CG) represents the collective set where malicious packages co-exist in the same security report. The security analysts share the findings in the report to reveal the attack campaign behind the malicious package. If two nodes have a co-existing edge , we put them into the same CG.
III-C Implementation & Validity
Implementation. We have implemented the prototype system of MalGraph via several open-source libraries. For the knowledge graph, the Neo4j [59] is used to store the knowledge graph. Given a package, we extract its code from the package to calculate its signature. The hashlib [60] package is used to calculate the SHA256 of the malware code. The Packj [61], an open-source tool to extract the SBOM of a package, extracts the AST and metadata information of the package. We use the Pandas library to process the dataset. The web crawler is implemented by the Scrapy framework [62]. For the similarity edge, we use the “text-embedding-3-large” as the pretraining model in OPENAI and leverage the scikit-learn library [63] to implement the K-Means algorithm.
@@ III-C Implementation & Validity
MalGraph Validity. There are several concerns about the validity of the subgraph. (1) False Positives. There are two causes bringing about false positives. First, the similarity calculation may be inaccurate if two packages use similar codes but belong to two different groups. There is no ground truth dataset to validate the similarity relationship. Second, the reports may contain malicious packages not belonging to the same attack campaign. In this case, we provide a manual inspection to filter out the false positives. Given a cluster or a report, we manually inspect its content to determine whether it is a false positive. (2) False Negatives. There are two causes for false negatives: unavailable packages and missing security reports. So far, our dataset (Table I) contains the largest number of OSS malicious packages. Our report dataset (Table III) contains the largest number of security reports. In future work, we will continue to find and collect new malicious packages and security reports to improve the MalGraph coverage.
B.K (4,953) | M. (1,346) | M.D (2,915) | G.A (179) | S.i (1,545) | T. (3,201) | D.D (1,397) | P. (7,311) | So. (664) | I.B (62) | |
B.K(4,953) | 50 | 1,348 | 102 | 502 | 14 | 79 | 385 | 0 | 20 | |
M.(1,346) | 50 | 310 | 0 | 128 | 68 | 0 | 23 | 0 | 2 | |
M.D(2,915) | 1,348 | 310 | 0 | 0 | 6 | 17 | 243 | 0 | 0 | |
G.A(179) | 102 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 2 | |
S.i(1,545) | 502 | 128 | 0 | 0 | 244 | 0 | 16 | 0 | 0 | |
T.(3,201) | 14 | 68 | 6 | 0 | 244 | 0 | 539 | 4 | 0 | |
D.D(1,397) | 79 | 0 | 17 | 0 | 0 | 0 | 12 | 0 | 0 | |
P.(7,311) | 385 | 23 | 243 | 0 | 16 | 539 | 12 | 0 | 0 | |
So.(664) | 0 | 0 | 0 | 0 | 0 | 4 | 0 | 0 | 0 | |
I.B(62) | 20 | 2 | 0 | 2 | 0 | 0 | 0 | 0 | 0 |
IV Insights of Malicious Package Group
@@
RQ1 (Redundancy & Quality): What impact does ad-hoc research have on malicious package datasets?
RQ2 (Diversity): How does the knowledge graph perform when applied to represent the diversity of malicious packages?
RQ3 (Attack Campaign): What attack campaign can the malicious packet group reflect?
RQ4 (Evolution): What is the evolution of malicious packages in an attack campaign over time?
IV-A Overlapping Degree (RQ1)
The research about the malicious package is done in an ad-hoc manner, indicating that cooperative relationships seem rare. We use the DG to represent duplicated malicious packages from different sources. Table IV shows the overlapping matrix of all sources. The left-upper part in the matrix represents the overlapping number between different sources from the academia. The right-lower part represents the overlapping number of different sources from the industry. The left-lower/right-upper part represents the overlapping number between academia and industry.
We observe that academia has a large overlapping degree (high values in left-upper/left-lower/right-upper parts), indicating a large redundancy. When malicious packages come from academia, they are more likely to be duplicated by other sources from academia or industry. The plausible reason is that academia does not detect malicious packages by itself but reuses the detection result from the industry. In contrast, the industry has a small overlapping degree (small values in the right-lower part), indicating a small redundancy. The largest overlapping degree is 539 between the Tianwen [23] and Phylum [25], the second largest degree is 244 between the Snyk.io [22] and Tianwen [23]. Other commercial sources have a small overlapping degree, including GitHub Advisory [24], DataDog [20], and Socket [26]. A plausible reason is that all security corporations claim they are the first to detect malicious packages by themselves, leading to one business concern about fake propaganda.
Furthermore, we leverage the coverage rate to represent duplicated malicious packages from different sources. Figure 4 shows the CDF of DG size among 3 OSS ecosystems: NPM, PyPI, and RubyGems. If a package is only reported by one source, its DG size is equal to 1. A large DG size indicates that many online sources report the malicious package. We observe that 80% of malicious packages are reported by only one source, and only 10% are reported by more than three sources.
Our Insight is that integrating and crawling malicious packages from all sources is necessary because the degree of overlap is low except for three sources from academia. The low overlapping degree also indicates a lack of industry-wide and industry-academia collaboration.
IV-B Dataset Quality (RQ1)
@@
We measure the dataset quality using two metrics: the update frequency and the missing rate ().
The updated frequency is an important indicator of dataset quality because of the quickly developed software supply chain in the software community. Table V detailed the updated frequency information of all sources. The open-source datasets [19, 16, 17, 20] from academia lack maintenance and update. The plausible reason is that researchers in academia lack enough motivation to maintain and update new malicious packages. Their update frequencies are rare, close to 0. In contrast, the industry community is more motivated to find new malicious packages than academia, and the newly found malicious packages are always updated with time. Overall, sources from the industry are more active and update more frequently than sources from academia.
Source | Last update date | Frequency |
Backstabber-Knife [19] | May 2020 | Never update |
Maloss [16] | Aug.2023 | one per 3 month |
Mal-PyPI dataset [17] | Aug.2023 | Never update |
DataDog [20] | May. 2023 | Never update |
GitHub Advisory [24] | Oct.2023 | one per 6 month |
Snyk.io [22] | Dec. 2023 | one per 2 month |
Tianwen [23] | Dec. 2023 | one per 2 month |
Phylum [25] | Nov. 2023 | one per 1 month |
Socket [26] | Dec. 2023 | one per 1 month |
The reflects the dataset’s quality in all sources. Many malicious packages are unavailable in our dataset, and we only record package names and versions. Table VI lists the missing rate of all sources, where the single reflects how many unavailable packages we rely on a single source, and the overall reflects how many unavailable packages we supplement packages from other sources. We observe that many sources from the industry have a large (e.g., Socket, Phylum, Snyk.io, and Tianwen) because they do not provide malicious packages to the public. A low indicates that the dataset (e.g., DataDog, Maloss, and Mal-PyPI) is available, and we can download malicious packages directly. Backstabber-Knife [19] has a large because it does not provide a dataset to the public, where we have sent dozens of application requests to the authors. The overall trend is similar to the single . In other words, an unavailable malicious package cannot be found from a different source.
We provide two reasons the malicious package is unavailable in the registry mirrors, as shown in Figure 5. (1) The malicious package has an early release time. When released too early, the mirrors completed synchronization with the root registry, and the package was removed. (2) The package’s persistent time is too short. Once an adversary releases the malicious package to the register, the administrator detects and removes it quickly. The sync gap between the root registry and the mirrors is larger than the persistent time of the malicious package.
Source | Missing Num.# (Total Num.#) | Single | All |
Backstabber-Knife [19] | 3,928 (4,953) | 79.31% | 76.5% |
Maloss [16] | 3 (1,346) | 0.22% | 0.22% |
@@
GitHub Advisory [24] | 166 (179) | 92.74% | 92.7% |
Snyk.io [22] | 1,161 (1,545) | 75.2% | 75.2% |
Tianwen [23] | 1,772 (3,201) | 55.4% | 55.4% |
DataDog [20] | 0 (1,397) | 0.0% | 0.0% |
Phylum [25] | 6,669 (7,311) | 91.2% | 91.1% |
Socket [26] | 664 (664) | 100% | 100% |
Blogs [27, 28, 29] | 59 (62) | 95.16% | 93.5% |
Total | 14,422 (23,573) | 64.14% |
Insights that we have learned are as follows. The missing rate of the malicious package is high even though we collect them from all sources. If a malicious package is missing, it is difficult to find it from other sources, which harms and impedes malware research in the community. The intrinsic reason is that the package managers remove the malicious package in time. The only way to solve the missing rate is industry-academia cooperation.
IV-C Malicious Package Diversity (RQ2)
As aforementioned, the large number of unique malicious packages does not imply diversity because many packages may be similar or belong to the same group. We use three types of groups (DeG, SG, CG) to represent the diversity of malicious packages.
Table VII shows the overall group diversity: the group number and average size. We find 76 SGs in NPM, 36 SGs in PyPI, and 4 SGs in RubyGems. The average size of the SG is close to 30, covering a large part of malicious packages in our dataset. The packages in the same group have a high similarity with each other. Note that there are 6,320 unique malicious packages in our dataset (Table II), grouping 116 SGs. Despite the sheer volume of malicious packages, many use similar code segments, indicating attackers reuse similar code bases to develop malicious packages. On the other hand, today’s security tools can easily detect most malicious packages because they use similar attack behaviors. Unknown/sophisticated attack behaviors have yet to emerge or be detected.
We find 11 DeGs in NPM, 1 DeGs in PyPI, and 0 DeGs in RubyGems. Typically, the average size of the DeG is close to 2, where one package acts as a dependency library and another relies on it. The key point of DeG is that attackers use benign software packages to camouflage or hide malicious packages to evade security detection tools. We observe that the DeG covers a small part of malicious packages in our dataset, indicating that dependency obfuscation is not the primary attack method in the OSS ecosystem. We find 26 CGs in PyPI, 50 CGs in NPM, and 6 CGs in RubyGems. Unlike the DeG and SG, the CG is obtained by manual industry analysts. The average size of the CG is close to 26, covering a large part of malicious packages in our dataset.
OSS | SG Num.# Ave. Size | DeG Num.# Ave. Size | CG Num.# Ave. Size |
NPM | 76 (17.78) | 11 (2.36) | 50 (46.1) |
PyPI | 36 (137.17) | 1 (2) | 26 (22.69) |
RubyGems | 4 (7.75) | 0 | 6 (7.67) |
Our Insight is that malicious package diversity in OSS ecosystems is not high due to the potential presence of similar behaviors in the code base. This finding indicates that today’s tools show a hallucination effect when detecting malicious packages because of similar attack behaviors rather than unknown/sophisticated ones.
IV-D Attack Campaign (RQ3)
Malicious packages cannot describe the context around how the malware is used and by whom. The life cycle of a malicious package is characterized into four distinct phases: development, release, detection, and removal, as shown in Figure 6. The first phase starts when an adversary or a hacker develops a malicious package. The next phase is when an adversary releases it into the OSS ecosystems, which is supposed to be done through deceptive tactics or a coordinated process. Then, a security researcher, a security company, or a package management administrator detects the malicious package. If the malicious package is first found by a security researcher or the security company, they will notify the package manager administrator. Finally, the package administrator removes the malicious package from the OSS ecosystem. The removal stage is not the end of the OSS malicious package. Adversaries have to change package operations to conduct subsequent attack campaigns. The reason is straightforward: the same package cannot be released again if it is removed from the registry. To conduct a subsequent attack, adversaries must change the malicious package, e.g., the package name, version, and source code. A malicious package is changed and re-released into the OSS ecosystems. Hence, the procedure {changingreleasedetectionremoval} will continue to cycle until the attacker stops the attack.
Similar Attack Campaign indicates that the same actor group releases several similar malicious packages to attack the OSS ecosystem. The operation is to change the package names or versions but not change the source code. Hence, we use the SG to represent similar attack campaigns. The average size of SG in PyPI is 137.17, and the average size of SG in NPM is 17.78 (Table VII). This finding indicates that a slight modification with the same malware is the primary attack method in the OSS ecosystem. The advantage of a similar attack campaign is that it is simple and has low costs. The disadvantage is that today’s security tools easily detect this attack campaign.
Dependency Attack Campaign indicates that the attacker releases a package with malicious codes and then creates a second one, depending on the first one. Figure 7 plots the attack based on the dependency libraries. The attacker first releases a package with malicious codes to the registry repository and then creates a second one, depending on the first one. The second package seems like a legitimate package without malicious behaviors. When the user downloads and installs the second package, the first package is automatically downloaded to the project, and the attack is triggered.
Complicated Attack Campaign indicates a series of malicious packages released by the same actor group, but we lack its context. Security reports record the context of the attack campaign, and the attacker uses those malicious packages to attack the OSS ecosystem. For example, the report (phylum.io) [64] claimed that the actor group released 15 malicious packages in the NPM ecosystems between August/9/2023 and August/19/2023. Figure 8 shows the release timeline of those malicious packages. Another example is the security report [65], where 5,943 malicious packages were released as the registering flood attack campaign towards the PyPI registry. The attacker released 5,943 malicious packages in a short time.
Active Period of Attack Campaign. Each group has a starting time to distribute its first malware and the last time to distribute its last malware. We use to denote the timestamp of the first malware and to denote the timestamp of the last malware. The time span between and represents the active period of the group. The active period is used to represent the time span of the attack campaign, where the malware is active intermittently in the OSS ecosystem.
Figure 9 plots the CDF of the active period of groups, including DeG, SG, and CG. 80% DeG attack campaigns have less than three years of active periods; 80% DeG attack campaigns have less than one year of active periods; and 80% SG attack campaigns have several days of active periods; Compared to SG and CG, DeG has the longest active period. This may be attributed to different attack strategies. DeG often first releases a benign package, allowing it to keep a long period before releasing a malicious package for an attack attempt. We observe that the SG attack campaign has the shortest active period, and it leverages a quick and simple attack strategy to deceive users into downloading and installing malicious packages. The CG attack campaign may constitute denial-of-service attacks, where attackers often simultaneously release a large number of malicious packages to achieve their objectives. We also find a long-tail effect of the active period for those attack campaigns, indicating some malicious packages have been released after more than ten years.
Insights that we have learned are as follows. (1) Although using a dependency library can hide malicious packages, the attacker does not use it. Simple and repeat attacks are commonly adopted in today’s SSC attacks. (2) Many malicious packages lack the context of the attack campaign, and the security report is the evidence to reveal the attack campaign behind the malicious packages.
IV-E Attack Campaign Evolution (RQ4)
When an adversary develops a malicious package, he will launch several attempts to attack OSS ecosystems. The adversary reuses their already-made code base to launch more of the same attacks against OSS ecosystems with a slight operation, differing from the prior detected and removed malicious package. Figure 10 depicts an attack campaign of a malicious package with multiple release attempts. A changing operation represents the difference between a prior package and a later one. The is the diff(, ), where is the prior malicious package and is the later malicious package. A release attempt of a malicious package has its download number in the OSS ecosystem, denoted as . Given a group, the set is denoted as {, , }, and we sort the group by the release time of the package , where . We can obtain the operation set {, , ,} and the set of download number {, , ,}
Download Evolution. We use the download number to represent the impact of the malicious package, where a large download number indicates a high impact on users in the OSS ecosystem. For instance, the PyPI ‘fallguy’ package has 345 download times by software developers after its first release. Figure 11 depicts the download number evolution of the malware groups, where the X-axis indicates the order in which the package is released per group, and the Y-axis indicates quartile values of the download number. The box plot extends from the first quartile (Q1) to the third quartile (Q3) of the data, with a point at the median. Due to the large amount of data, we plotted the box for every ten X-axis labels. It is evident that the majority of release attempts exhibit a low download number (0 or 1). The reason is that the OSS registry detects malicious packages quickly and removes them in real-time. A minority of release attempts demonstrate a large download number (1040). Furthermore, we find several release attempts of malicious packages have millions of download numbers, denoted as outliers. The reason is that those outliers belong to popular packages where one version is denoted as the malware. Therefore, their download numbers are equal to the ones for the latest version of popular packages. Although outlier points are only a small part of our dataset, they have a great impact on the distribution of the box plot in Figure 11.

### v3

@@ III-D Co-existing Edge
Co-existing Edge indicates that two malicious packages co-exist in the same security report. Malicious packages cannot describe the context around how the malware is used and by whom. Hence, the security report is the evidence that reveals the attack campaign behind the malicious package. For example, the security report [59] claimed that an adversary author ‘Lolip0p’ developed malicious packages called ‘Colorslib’, ‘httpslib’ and ‘libhttps’, where the co-existing relationship ‘Colorslib’, ‘httpslib’, ‘Colorslib’, ‘libhttps’, and ‘httpslib’, ‘libhttps’. Another example is the security report [60], where 5,943 malicious packages were released as a register flood attack campaign towards the PyPI registry. The attacker released 5,943 malicious packages in a short time.
We determine a co-existing relationship between two malicious packages if they co-exist in the same security report. A dataset with the security report is essential for building the co-existing edge in the malicious package graph. However, there is no public dataset with the security report.
@@
To address this practical challenge, we propose a web crawler to collect the security analysis reports from the Internet. The web crawler is implemented by the Scrapy framework [61]. (1) We use the commercial websites and individual blogs (Section II-B) as the initial seed for finding the security analysis reports. (2) We leverage the search engine (Google) to find similar URLs to the initial URL. (3) We use the web crawler to collect the security analysis reports from the Internet. (4) We manually filter out irrelevant web pages and keep the security reports about malicious packages. Through the manual approach, we have found 1,366 reports that provide security analysis about malicious packages. Table III lists details of the online websites and security reports. Our dataset contains the largest number of security reports about OSS malicious packages.
@@
IV Dataset and MalGraph Validity
There is no established ground truth for OSS malware or their relationships. As a result, it is impossible to provide a systematic evaluation of the data collection approach (Section II-B) or MalGraph (Section III). Instead, we conducted controlled experiments and performed manual inspections to guarantee the reproducibility of our results.
IV-A Malware Dataset Validity
Malware Correctness. (1) Manual Inspection. We assume that online sources are reliable and have a good reputation in the security community. If a legitimate package is falsely reported as a malicious package, it may be added to contain our dataset. Due to the large number of malicious packages, manually inspecting each package via the reserved engineering technique is impractical. We use a simple rule to remove the false positives: if the root register does not remove a package, it is not a malicious package. (2) Controlled experiment. We conducted five sampling experiments from our collected dataset. In each experiment, we randomly selected 100 malicious packages. We utilized Semgrep [62] and GuardDog [63], two commercial state-of-the-art tools, to scan the sampled packages. Although these tools may produce a high number of false positives, we manually inspected the scanning results to verify whether the packages were indeed malicious. The experiment demonstrated that 100% sampling packages belong to OSS malware.
Malware Dataset Transparency. To guarantee the reproducibility of analysis results, we provide the malware dataset to the community. Due to page limitations, we present only a summary of datasets in the paper, while the detailed dataset is available in the GitHub repository. Note that malicious packages are not available in the GitHub repository because of ethical considerations, i.e., script kiddies. In contrast, we share all malicious packages (Table I) via a private GitHub repo, and applicants need to send an email request to join this repo. We believe that a high-quality and complete OSS malware dataset is an important source for our community. With it, many researchers can avoid repeating such challenging and tedious work in the SSC field.
IV-B MalGraph Validity
@@
MalGraph Correctness. There are several concerns about the validity of the MalGraph. (1) False Positives. There are two causes bringing about false positives. First, the similarity calculation may be inaccurate if two packages use similar codes but belong to two different groups. There is no ground truth dataset to validate the similarity relationship. Second, the reports may contain malicious packages not belonging to the same attack campaign. In this case, we provide a manual inspection to filter out the false positives. Given a cluster or a report, we manually inspect its content to determine whether it is a false positive. (2) False Negatives. There are two causes for false negatives: unavailable packages and missing security reports. So far, our malware dataset (Table I) contains the largest number of OSS malicious packages. Our report dataset (Table III) contains the largest number of security reports. In future work, we will continue to find and collect new malicious packages and security reports to improve the MalGraph coverage.
MalGraph Transparency. To guarantee the reproducibility of analysis results, we provide the security report dataset and the knowledge graph to the community. We list all package groups (manual labeling) so researchers can identify which package to use in the dataset. Our analysis contains all malicious packages from Table I.
B.K (5,937) | M. (1,223) | M.D (2,915) | G.A (179) | S.i (1,540) | T. (3,151) | D.D (1,387) | P. (7,299) | So. (664) | I.B (62) | |
B.K(5,937) | 368 | 2,897 | 0 | 3 | 36 | 7 | 966 | 1 | 36 | |
M.(1,223) | 368 | 201 | 6 | 1 | 69 | 0 | 0 | 3 | 6 | |
M.D(2,915) | 2,897 | 201 | 0 | 0 | 32 | 7 | 918 | 0 | 0 | |
G.A(179) | 0 | 6 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | |
S.i(1,540) | 3 | 1 | 0 | 0 | 106 | 0 | 17 | 0 | 0 | |
T.(3,151) | 36 | 69 | 32 | 0 | 106 | 0 | 272 | 2 | 0 | |
D.D(1,387) | 7 | 0 | 7 | 0 | 0 | 0 | 15 | 0 | 1 | |
P.(7,299) | 966 | 0 | 918 | 0 | 17 | 272 | 15 | 0 | 0 | |
So.(664) | 1 | 3 | 0 | 0 | 0 | 2 | 0 | 0 | 0 | |
I.B(62) | 36 | 6 | 0 | 1 | 0 | 0 | 1 | 0 | 0 |
V Insights of Malicious Package Group
@@
RQ1 (Quality): What impact does ad-hoc research have on malicious package datasets?
RQ2 (Diversity): How is the diversity of the OSS malicious package dataset?
RQ3 (Dependent-hidden): Would dependent-hidden malware bring a great attack advantage?
RQ4 (Malware Context): What is the context around the OSS malware packages?
V-A Dataset Quality (RQ1)
@@
We measure the dataset quality using the overlapping degree and the missing rate ().
Overlapping. The research about the malicious package is done in an ad-hoc manner, indicating that cooperative relationships seem rare. We use the duplicated edge to represent the overlapping degree of different sources: if two nodes have a duplicated edge , we put them () into the same subgraph. Table IV shows the overlapping matrix of all sources. The left-upper part in the matrix represents the overlapping number between different sources from the academia. The right-lower part represents the overlapping number of different sources from the industry. The left-lower/right-upper part represents the overlapping number between academia and industry.
We observe that academia has a large overlapping degree (high values in left-upper/left-lower/right-upper parts), indicating a large redundancy. When malicious packages come from academia, they are more likely to be overlapped with other sources from academia or industry. Obviously, Backstabber-Knife [18] (B.K) has the largest overlapping degree with other sources (respectively, 368 with M., 2,897 with M.D, 3 with S.i, 36 with T., 7 with D.D, 966 with P., 1 with So., 36 with I.B). It is because Backstabber-Knife [18] is the source that integrates other datasets into its dataset. In contrast, the industry has a small overlapping degree (small values in the right-lower part), indicating a small redundancy. The largest overlapping degree is 272 between the Tianwen [22] and Phylum [24], the second largest degree is 106 between the Snyk.io [21] and Tianwen [22]. Other commercial sources have a small overlapping degree, including GitHub Advisory [23], DataDog [19], and Socket [25]. The plausible reasons are as follows: (1) academia does not detect malicious packages by itself but reuses the detection result from the industry; (2) in the industry, all security corporations claim they are the first to detect malicious packages by themselves, leading to small overlapping degree.
Figure 6 shows the CDF of the number of occurrences of a malicious package in all online sources. If a package is reported by only one source, its occurrence count is equal to 1; otherwise, it is greater than or equal to 2. We observed that the number of occurrences for all packages is less than or equal to 4. In NPM and RubyGems ecosystems, 80-90% of malicious packages appear one time and almost no package appears more than 3 times. Approximately 50% of the packages on PyPI appear only once. Of the remaining, 25% occur twice, another 25% occur three times, while only 56 packages are observed to appear four times.
In the following analysis (Missing Rate, RQ2, RQ3, and RQ4), the dataset we used has removed the overlap.
Source | Missing Num.# (Total Num.#) | Local | Global |
Backstabber-Knife [18] | 0 (5,937) | 0.00% | 0.00% |
Maloss [16] | 0 (1,223) | 0.00% | 0.00% |
@@
GitHub Advisory [23] | 166 (179) | 92.74% | 89.39% |
Snyk.io [21] | 1,160 (1,540) | 75.32% | 75.26% |
Tianwen [22] | 1,769 (3,151) | 56.14% | 53.70% |
DataDog [19] | 0 (1,387) | 0.00% | 0.00% |
Phylum [24] | 6,606 (7,298) | 90.52% | 80.49% |
Socket [25] | 664 (664) | 100% | 99.40% |
Blogs [26, 27, 28] | 59 (62) | 95.16% | 30.65% |
Total | 10,424 (24,356) | 39.27% |
Missing Rate. The reflects the dataset’s quality in all sources because many malicious packages are unavailable in our dataset. To measure the dataset quality, we propose two metrics: local and global missing rate (). Local is equal to , where is the number of unavailable packages in the ith source, and is the number of packages in the ith source. Global is calculated as follows,
where is missing in the ith source, and is the set of packages from other sources. The is an indicator symbol: if other sources can supplement the missing package, it is assigned to 0; otherwise, it is assigned to 1. In other words, local reflects unavailable packages that rely on a single source, and global reflects package unavailable from all sources.
Table V lists missing rates of all sources. Local in the open-source dataset (e.g., DataDog, Maloss, and Mal-PyPI) is low because we can download malicious packages directly from those sources. Backstabber-Knife is an important source to supplement the missing packages because it integrates malicious packages from different sources. Global in ‘Blogs’ source is significantly lower than its local because the open-source datasets supplement those missing packages. We observe that many sources from the industry have a large (e.g., Socket, Phylum, Snyk.io, and Tianwen) because they do not provide malicious packages to the public. Meanwhile, those sources have similar Global and Local , indicating that an unavailable malicious package from those sources cannot be supplemented by a different source.
We further observe the release time of unavailable malicious packages. We query the release time to the official registry via the specific package name and version. Specifically, there are 24,356 malicious packages and 10,424 missing packages in our dataset. Figure 7 shows the release time of packages from 2014 to 2024, with the X-axis showing the release time and the Y-axis showing the package number. The blue bar represents unavailable packages, the white bar indicates all packages. The time of missing packages covers a span of 6 years. We find that the number of the missing package has a peak in February 2023. It is the reason that 5,235 malicious packages acted as the registration flooding attack against the PyPI ecosystem. There are two reasons malicious packages were unavailable in the registry mirrors, as shown in Figure 8. (1) The malicious package has an early release time. When released too early, the mirrors completed synchronization with the root registry, and the package was removed. Figure 7 shows the release time of missing malware packages across all sources, many of which were released 3-4 years ago. (2) The package’s persistent time is too short. Once a malware author releases a malicious package to the register, the administrator detects and removes it quickly. The sync gap between the root registry and the mirrors is larger than the persistent time of the malicious package.
OSS | Pkg. Num.# | Subgraph Num.# | Ave. Num.# | Largest Size |
NPM | 2,994 | 157 | 19.07 | 827 |
PyPI | 4,365 | 295 | 14.80 | 829 |
RubyGems | 83 | 37 | 2.24 | 6 |
V-B Malware Diversity (RQ2)
A large number of malicious packages does not imply diversity if many packages are similar. We use similar code bases to represent OSS malware diversity: if two nodes have a similar edge , we put them into the same subgraph.
Table VI lists the overall similar subgraphs based on similar code bases. We find 157 subgraphs in NPM, 295 subgraphs in PyPI, and 37 subgraphs in RubyGems. The subgraph average size indicates the number of malicious packages with similar code bases. The average number of subgraphs in PyPI contains 14.80 malicious packages; the average size in NPM is 19.07; and the subgraph in RubyGems has 2.24 packages on average. The largest size in NPM is 827 packages, and the one in PyPI has 829 packages. We further inspected malicious code in the largest PyPI subgraph (829). All packages shared the same payload that was obfuscated with variable names in Chinese. After debugging, we found that this malware changed shortcuts to load browsers extensions. The browser extension replaced Crypto wallet addresses with hard-coded wallet addresses from attackers.
Despite the sheer volume of malicious packages, low software diversity brings about two implications. First, directly using the OSS malware dataset would lead to the bias of detection models and algorithms. If many malicious packages are centered around a few ones, the dataset is not representative, and the aggressive malicious packages dominate the dataset. Prior works [16, 64, 17] didn’t take into account the diversity of the malicious packages. Second, there is a need to define malware families for OSS malware. Traditional malware families are designed to illustrate the diversity of binary samples, which is not suitable for OSS malware.